A newspaper printer manager requires the proportion of errors to be no more than 4%. A random sample of 100 items results in 6 errors. Use a 5% significance level to test the claim that production has more than 4% errors.
Answer #1

Null hypothesis:

Alternative hypothesis:

Sample proportion = 6/100 = 0.06

n = 100

We will use z-test because we are dealing with proportion.

This is one-tail test. So, p-value = P(z > 1.02) = 0.1539 (from z-table)

As p-value is greater than significance level.

So, we fail to reject null hypothesis.

So, at 5% level of significance (or 95% confidence interval) we don't have sufficient evidences to support the claim that the production has more than 4% errors.

Please comment if any doubt. Thank you.